PRECISION SPECTROSCOPY OF TELLURIUM

J. COKER, J.E. FURNEAUX, Department of Physics and Astronomy, University of Oklahoma, Norman, OK 73069.

Tellurium (Te_{2}) is widely used as a frequency reference, largely due to the fact that it has an optical transition roughly every 2-3 GHz throughout a large portion of the visible spectrum. Although a standard atlas encompassing over 5200 cm^{−1} already exists [1], Doppler broadening present in that work buries a significant portion of the features [2]. More recent studies of Te_{2} exist which do not exhibit Doppler broadening, such as Refs. [3-5], and each covers different parts of the spectrum. This work adds to that knowledge a few hundred transitions in the vicinity of 444 nm, measured with high precision in order to improve measurement of the spectroscopic constants of Te_{2}’s excited states. Using a Fabry Perot cavity in a shock-absorbing, temperature and pressure regulated chamber, locked to a Zeeman stabilized HeNe laser, we measure changes in frequency of our diode laser to ∼1 MHz precision. This diode laser is scanned over 1000 GHz for use in a saturated-absorption spectroscopy cell filled with Te_{2} vapor. Details of the cavity and its short and long-term stability are discussed, as well as spectroscopic properties of Te_{2}. References:

1. J. Cariou, and P. Luc, Atlas du spectre d’absorption de la molecule de tellure, Laboratoire Aime-Cotton (1980).

2. J. Coker et al., J. Opt. Soc. Am. B 28, 2934 (2011).

3. J. Verges et al., Physica Scripta 25, 338 (1982).

4. Ph. Courteille et al., Appl. Phys. B 59, 187 (1994)

5. T.J. Scholl et al., J. Opt. Soc. Am. B 22, 1128 (2005).